OpenStudy (idku):

I am stuck on this baseball chart for 5 days. All I need is to answer some 3 questions. Can someone just help me with the right formulas?

OpenStudy (idku):

Michael Jordan is widely known for his legendary basketball career. He is a said to be one of the greatest athletes of all time. In 1993, Jordan retired from the NBA to pursue a professional baseball career. Do you think his baseball career damaged his renowned athletic career? The following chart contains Michael Jordan's Baseball Stats. Research baseball statistics online to determine what each labels represent in the baseball world. Year Avg Gm AB R H TB 2B 3B HR 1994 0.202 127 436 46 88 116 17 1 3 1) How can you use the given information to determine how many singles Jordan hit during his baseball career? Do the calculation and show your work. 2) Find MJ's slugging average. Do the calculation and show your work. 3) Based on his slugging average, do you think Jordan was a good baseball player? Do you think baseball damaged his renowned athletic career? ──────────────────────────────────────

OpenStudy (idku):

@ParthKohli @mathmale, @ganeshie8 @No.name

ganeshie8 (ganeshie8):

http://en.wikipedia.org/wiki/Baseball_statistics

ganeshie8 (ganeshie8):

http://prntscr.com/3t12ha

OpenStudy (idku):

Can someone just tell me the correct formula for number 1 ? I am very tired and I was stuck on this since ever-:(

ganeshie8 (ganeshie8):

I think your table has a typo : Year Avg Gm AB R H \(\large \color{Red}{1B}\) 2B 3B HR 1994 0.202 127 436 46 88 116 17 1 3

OpenStudy (idku):

I'll check

OpenStudy (idku):

It says TB in my lesson. I don't know why, are you sure it is 1B and not TB ?

ganeshie8 (ganeshie8):

I have no clue about baseball - I thought it could be a typo based on the link I gave you earlier >.<

ganeshie8 (ganeshie8):

In that link it says \(1B = Singles\) but in your table there is no such thing as \(1B\), so I am really not not that sure

OpenStudy (idku):

They wouldn't ask me to find the singles if the were already given

ganeshie8 (ganeshie8):

thats a good point :) lets do some more research

OpenStudy (idku):

Yeah, I would think that my table s correct then.

OpenStudy (idku):

So how do I find the singles ?

ganeshie8 (ganeshie8):

First, we need to understand what TB represents

OpenStudy (idku):

Total Bases

ganeshie8 (ganeshie8):

yeah, hows is it calculated ?

ganeshie8 (ganeshie8):

`TB – Total bases: one for each single, two for each double, three for each triple, and four for each home run [H + 2B + (2 * 3B) + (3 * HR)] or [1B + (2 * 2B) + (3 * 3B) + (4 * HR)] `

ganeshie8 (ganeshie8):

Year Avg Gm AB R H TB 2B 3B HR 1994 0.202 127 436 46 88 116 17 1 3

ganeshie8 (ganeshie8):

use the last four columns : TB = 1*singles + 2*doubles + 3*triples + 4*quadruples

ganeshie8 (ganeshie8):

116 = 1*singles + 2*17 + 3*1 + 4*3 solve `singles`

OpenStudy (idku):

Okay, I'll solve for singles:)

OpenStudy (idku):

116 = 1×(1B) + 2×17 + 3×1 + 4×3 116 = 1×(1B) + 34 + 3 + 12 116 = 1B + 49 67 = 1B

OpenStudy (idku):

Okay and how do I find the slugging average ?

ganeshie8 (ganeshie8):

Looks good to me http://www.wolframalpha.com/input/?i=116-2*17-3*1-4*3

ganeshie8 (ganeshie8):

search for `SLG` in the earlier wiki link, looks it also has some formula...

ganeshie8 (ganeshie8):

http://en.wikipedia.org/wiki/Slugging_average

ganeshie8 (ganeshie8):

SLG = TB/AB

OpenStudy (idku):

SLG = [ 1×(1B) + 2×(2B) + 3×(3B) + 4×(HR) ] / AB SLG = [ 1×(67) + 2×(17) + 3×(1) + 4×(3) ] / 436 SLG = [ 67 + 34 + 3 + 12 ] / 436 SLG = [ 116 ] / 436 SLG = 116 / 436 ≈ 0.27 right ?

ganeshie8 (ganeshie8):

yes, but TB was already given in the table right ?

OpenStudy (idku):

Yes. So either way I am correct. Right ?

ganeshie8 (ganeshie8):

Yep !

OpenStudy (idku):

So he was therefore not a good player, if his slugging average is 0.27

OpenStudy (idku):

(question 3 )

ganeshie8 (ganeshie8):

not sure, @nincompoop

OpenStudy (idku):

Yeah, I just really know almost nothing about baseball -:(